OpenStudy (anonymous):

Could I get some help with this? Solve √x=4 for x

Nnesha (nnesha):

there is square root of x to cancel out square root you have to take square both side

TheSmartOne (thesmartone):

\(\bf\huge~~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @Taylorcheer98

Nnesha (nnesha):

\[\huge\rm \sqrt{x} = x^\frac{ 1 }{ 2 }\] so when u take square root square root cancel out

OpenStudy (anonymous):

Would my answer be 4?

TheSmartOne (thesmartone):

Is your question? \(\Large\sqrt{x}=4\)

OpenStudy (academicgurusinc):

You have first square each side of the equation. When you do that, you'll get x = 16 as your final answer. Hope that helps!

Nnesha (nnesha):

nope take square both side

OpenStudy (anonymous):

Yes

Nnesha (nnesha):



OpenStudy (anonymous):

The square root so 16 is 4, so 4?

Nnesha (nnesha):

not square root take square

OpenStudy (anonymous):

Or 16?

Nnesha (nnesha):

what is \[\huge\rm (\sqrt{x})^2\] = ????

OpenStudy (anonymous):

I'm not sure how I do that

TheSmartOne (thesmartone):

remember that \(\LARGE\sqrt{x}=x^\frac{1}{2}\)

Nnesha (nnesha):

 like i said before square root of x can be written as 1/2 exponent so what is \[\huge\rm (x^{\frac{ 1 }{ 2 }})^2\]